y= -x+2 y= -2x+6
i have to answer in ordered pair. please help.i am sorry but i have no textbook as all work is done in an apex form on internet.

Answer by checkley71(8403)

Y=-X+2
Y=-2X+6
-X+2=-2X+6
-X+2X=6+2
X=8 ANSWER.
Y=-8-2
Y=-10 ANSWER.
PROOF:
10=-2*8+6
10=-16+6
10=10